OpenStudy (anonymous):

A 10.0 Ω resistor is in parallel with a 30.0 Ω resistor and supplied by a 45.0 V battery. What is the total current drawn from the battery? A. 0.003 A B. 1.10 A C. 6.00 A D. 338 A

OpenStudy (michele_laino):

here the total resistance is: \[\Large R = \frac{1}{{\frac{1}{{10}} + \frac{1}{{30}}}} = ...\]

OpenStudy (anonymous):

7.5?

OpenStudy (michele_laino):

ok! so the requested current is: \[\Large I = \frac{{\Delta V}}{R} = \frac{{45}}{{7.5}} = ...amperes\]

OpenStudy (anonymous):

6! so our solution is choice C? 6.00 A?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay! thanks!:D

OpenStudy (michele_laino):

:)